# Spatial Policy: Guiding Visuomotor Robotic Manipulation with Spatial-Aware Modeling and Reasoning

### v1

@@ Abstract
Vision-centric hierarchical embodied models have demonstrated strong potential for long-horizon robotic control. However, existing methods lack spatial awareness capabilities, limiting their effectiveness in bridging visual plans to actionable control in complex environments. To address this problem, we propose Spatial Policy (SP), a unified spatial-aware visuomotor robotic manipulation framework via explicit spatial modeling and reasoning. Specifically, we first design a spatial-conditioned embodied video generation module to model spatially guided predictions through a spatial plan table. Then, we propose a spatial-based action prediction module to infer executable actions with coordination. Finally, we propose a spatial reasoning feedback policy to refine the spatial plan table via dual-stage replanning. Extensive experiments show that SP significantly outperforms state-of-the-art baselines, achieving a 33.0% average improvement over the best baseline. With an 86.7% average success rate across 11 diverse tasks, SP substantially enhances the practicality of embodied models for robotic control applications. Code and checkpoints are maintained at https://plantpotatoonmoon.github.io/SpatialPolicy/.
Introduction
In recent years, hierarchical embodied models combining high-level planning and low-level action have achieved remarkable success in long-horizon robotic manipulation (Brohan et al. 2024; Cheang et al. 2024; Ze et al. 2024; Li et al. 2025b). Among these approaches, vision-based methods, known as visuomotor robotic manipulation, have emerged as a particularly promising direction. By leveraging rich visual cues, these methods enable effective cross-hierarchical information transfer across diverse robotic embodiments and environments (Du et al. 2023; Tian et al. 2024; Ko et al. 2023; Soni et al. 2025; Black et al. 2023; Li et al. 2025a). Specicially, visuomotor robotic manipulation employs a generative model to predict trajectories of future video, which are subsequently translated into executable actions through a action prediction module.
Despite significant advances, existing visuomotor robotic manipulation frameworks lack spatial awareness ability, failing to effectively bridge visual plans to executable actions in complex environments (Tian et al. 2024; Black et al. 2023). For instance, AVDC (Ko et al. 2023) lacks explicit spatial modeling in its video generation process, resulting in physically implausible predictions, such as robotic arms passing through walls. Feedback module of CLOVER (Bu et al. 2024) enhances dynamic environment adaptation, but its inability to process spatial relationships prevents handling of occluded target scenarios (Tian et al. 2024).
@@ Introduction
To address these challenges, we propose Spatial Policy (SP), a unified framework for robotic manipulation that integrates spatial-aware video generation, action prediction, and feedback. The core idea is to unify spatial modeling and reasoning via a spatial plan table that guides video generation, enables spatially consistent action prediction, and supports feedback-driven refinement during execution. Specifically, (1) we propose a Spatial-Conditioned Embodied Video Generation module that generates visual predictions guided by a structured spatial plan table, which encodes object-goal relations and directional cues to ensure spatial coherence in imagined futures, addressing the challenge of modeling spatially meaningful video trajectories; (2) we introduce a Spatial-Based Action Prediction module that infers executable actions by capturing spatial motion dynamics between predicted frames using spatial coordinates, tackling the challenge of translating visual predictions into spatially consistent actions; (3) we develop a Spatial Reasoning Feedback Policy that refines the spatial plan and video execution through dual-stage replanning based on VLM feedback and rule-based corrections, solving the challenge of real-time spatial reasoning during policy execution. Extensive experiments demonstrate that our model outperforms existing baselines, achieving 86.7% performance. The contributions of this paper are summarized as follows:
@@
  • Modular design with structured spatial modeling and reasoning. SP consists of three synergistic modules: a spatial-conditioned video generation module guided by a structured spatial plan table, a spatial-based action prediction that preserves spatial consistency, and a spatial reasoning feedback policy that refines execution through real-time spatial replanning.
@@
  • Comprehensive evaluation and strong performance. We conduct extensive experiments on the MetaWorld benchmark to assess the impact of spatial integration, including ablations on video generation, action prediction, and feedback mechanisms. SP outperforms prior methods by over 30%, demonstrating the effectiveness of spatial reasoning in visuomotor robotic manipulation.
Related Work
Video Generation for Embodied Manipulation
@@
Video generation has recently been explored as a visual planning mechanism for embodied agents. AVDC (Ko et al. 2023) generates action-annotated video plans from raw RGB inputs using dense correspondencesenabling policy learning without action supervision. However, due to the lack of explicit spatial reasoning in both generation and action prediction, limiting precision in manipulation tasks. VideoAgent (Soni et al. 2025) refines video plans via self-conditioning to reduce hallucinations and improve visual quality. However, it inherits the lack of explicit spatial reasoning from AVDC in both generation and action prediction, limiting precision in geometry-sensitive manipulation tasks. Clover (Bu et al. 2024) introduces a closed-loop framework combining video diffusion planning and feedback control via an embedding-based error signal, achieving strong results in long-horizon tasks. However, its spatial modeling is limited to depth-based cues and lacks explicit 3D spatial planning or structured geometric reasoning, which restricts generalization across diverse spatial layout. SEER (Tian et al. 2024)proposes a Predictive Inverse Dynamics Model (PIDM) that forecasts future visual states and infers actions end-to-end, closing the loop between vision and control. While it achieves strong performance through large-scale training, it lacks explicit spatial reasoning modules and does not leverage structured spatial plans, which may limit interpretability and adaptability in geometry-sensitive tasks. SuSIE (Black et al. 2023) leverages an image-editing diffusion model to generate visual subgoals for a low-level goal-conditioned policy, enabling better generalization in unstructured environments. However, its subgoal generation relies on appearance-based edits without explicit incorporation of 3D spatial relationships, limiting its precision in spatially complex manipulation scenarios.
Spatial Reasoning in Vision-Language Embodied Systems
@@
Spatial understanding has been addressed in embodied AI from several perspectives. SpatialVLM (Chen et al. 2024a) and GraspCorrect (Lee, Hong, and Kim 2025) leverage pretrained vision-language models to extract chain-of-thought spatial cues and determine object relations in manipulation tasks. 3D-VLA (Zhen et al. 2024) explicitly encodes spatial semantics by aligning 3D vision with textual instructions. SPA (Zhu et al. 2025) utilizes point cloud reconstruction for spatial alignment in manipulation and navigation tasks. Meanwhile, LLM-Planner (Song et al. 2023)and related methods employ large language models to infer spatially structured plans in symbolic form.
Despite these advances, such methods either focus on semantic spatial understanding or operate in symbolic domains, lacking visual feedback and executable rollout. Moreover, most rely on handcrafted prompts or static spatial priors and do not support continuous video generation. In contrast, our work closes this gap by embedding structured spatial plans—automatically derived from environment geometry—into a generative video prediction framework, thereby tightly coupling spatial reasoning and visual synthesis for action-aware planning.
Spatial Conditioning in Generative Video Models
Recent advances in text-to-video diffusion models, such as AnimateDiff (Guo et al. 2024) and VideoCrafter2 (Chen et al. 2024b), achieve high-fidelity generation but lack controllability over agent behavior or scene dynamics. ScenarioDiff (Zhang et al. 2025) and DirecT2V (Hong et al. 2024) propose spatially guided generation using manually specified constraints like bounding boxes or motion paths. These constraints are often synthetic and decoupled from real embodied agent states, limiting their adaptability and semantic alignment with embodied control tasks.
Spatial Policy
We propose a visuomotor robotic manipulation framework that integrates structured spatial reasoning into the perception-to-control pipeline, enabling robots to generate spatially consistent visual plans and execute them reliably. As illustrated in Fig. 1, our system comprises three interconnected modules. First, the Spatial-Conditioned Embodied Video Generation module generates future video trajectories conditioned on a structured spatial plan table composed of atomic actions, directional vectors, and relative distances. Second, the Spatial-Based Action Prediction module predicts actions aligned with the generated video using the spatial coordinates. Third, the Spatial Reasoning Feedback Policy monitors spatial alignment during execution and dynamically refines both the spatial plan table and the visual predictions. Together, these modules form a closed-loop framework that maintains spatial consistency across video imagination, action prediction, and feedback-based reasoning.
@@
Figure 1: Framework Overview of Spatial Policy. Our system comprises three modules: (1) Spatial-Conditioned Embodied Video Generation, which uses a Spatial Plan Table obtained via VLM reasoning over the spatial offset between the robot and the object, to guide diffusion model in generating spatially coherent video prediction. (2) Spatial-Based Action Prediction, which converts the generated video into executable actions using spatial coordinates. (3) Spatial Reasoning Feedback Policy, which enables real-time correction via a dual-stage replanning strategy, combining VLM-based video judgement and policy diagnostics to refine the spatial plan table for closed-loop control.
Spatial Conditioned Embodied Video Generation
@@
To generate spatially grounded visual trajectories for visuomotor robotic manipulation, we adopt a conditional diffusion framework for future video prediction. The model takes as input an initial observation frame and a textual task description, and generates a sequence of future frames representing high-level visual plans. To enhance spatial awareness, we additionally introduce spatial plan table derived from spatial relationships as an auxiliary global condition.
@@
To capture precise spatial intent, we extract the coordinates of the robot end-effector $\mathbf{p}_{\text{ee}}\in\mathbb{R}^{3}$ and the target object $\mathbf{p}_{\text{obj}}\in\mathbb{R}^{3}$ from the simulator, and compute their relative offset as $\Delta\mathbf{p}=\mathbf{p}_{\text{obj}}-\mathbf{p}_{\text{ee}}$. This spatial offset is passed into a pretrained visual-language model (GPT-4o) to produce a structured spatial plan. The generated plan consists of sequential subgoals, each encoding an atomic action with an action type, a direction vector, and a distance scalar, as illustrated in Fig. 1.
@@
All subgoal embeddings $\{\mathbf{e}_{1},\dots,\mathbf{e}_{n}\}$ are concatenated with the task text embedding $\mathbf{z}_{\text{task}}\in\mathbb{R}^{d}$ obtained via CLIP (Radford et al. 2021) to form the composite global condition:
@@
We follow the design of a u-net-based denoising diffusion probabilistic model, adapted from the Imagen architecture (Ho et al. 2022), consistent with AVDC (Ko et al. 2023). Given an initial frame $\mathbf{I}_{0}\in\mathbb{R}^{H\times W\times 3}$ and a composite condition vector $\mathcal{T}$ that incorporates both task instructions and structured spatial information, the model predicts a sequence of $T=7$ future frames $\mathbf{I}_{1:T}$, resulting in an 8-frame trajectory. The generative objective is to approximate the posterior distribution:
@@
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is sampled noise, $\beta_{t}$ is the diffusion schedule, and $\epsilon_{\theta}$ is the model’s denoising network. Further training details are provided in the Appendix.
Spatial Based Action Prediction
To convert spatially grounded visual plans into executable actions, we design an Spatial Coordinate-Guided Action Prediction module built upon the diffusion policy framework (Chi et al. 2023; Ji et al. 2025), due to its strong capacity for modeling multimodal distributions and complex temporal dependencies, making it particularly well-suited for capturing the spatial variability and uncertainty inherent in visual planning. The model predicts multi-step actions that guide the robot from the current observation toward the visual goal implicitly defined by the generated video sequence, leveraging spatial coordinate to capture fine-grained spatiotemporal dynamics.
@@
We condition our policy on three types of information: (1) the current image $\mathbf{I}_{\text{cur}}$, (2) the goal image $\mathbf{I}_{\text{goal}}$ (sampled from the generated visual plan within the next $k$ steps, i.e., $\mathbf{I}_{\text{goal}}=\mathbf{I}_{t+\delta}$ where $1\leq\delta\leq k$), and (3) the spatial coordinate of robotic arms $\mathbf{p}_{\text{cur}}\in\mathbb{R}^{3}$. These inputs capture both spatial and temporal motion cues. All inputs are encoded into compact feature vectors: the image pair is stacked along the channel dimension and passed through a shared ResNet encoder, while the spatial coordinate is encoded via a shallow MLP. The resulting modality-specific features are then concatenated to form a unified condition vector, which guides the denoising network in our diffusion policy-based action prediction model.
Diffusion Policy-based Action Prediction Model
We formulate action prediction as a conditional generative process. Specifically, we adopt a unet-based diffusion policy architecture to model the conditional distribution over action trajectories:
$p_{\theta}(\mathbf{a}_{0:H-1}\mid\mathbf{I}_{\text{cur}},\mathbf{I}_{\text{goal}},\mathbf{p}_{\text{cur}})$ (5)
@@
where $\mathbf{c}$ denotes the concatenated global condition vector from all encoded inputs, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise, and $t$ is the diffusion timestep. $\epsilon_{\theta}$ is the learned denoising network. At inference time, the model predicts the entire action trajectory in a denoising pass, and the resulting sequence is executed step-by-step in the environment. Further training details are provided in the Appendix.
Spatial Reasoning Feedback Policy

### v2

@@ Abstract
Vision-centric hierarchical embodied models have demonstrated strong potential. However, existing methods lack spatial awareness capabilities, limiting their effectiveness in bridging visual plans to actionable control in complex environments. To address this problem, we propose Spatial Policy (SP), a unified spatial-aware visuomotor robotic manipulation framework via explicit spatial modeling and reasoning. Specifically, we first design a spatial-conditioned embodied video generation module to model spatially guided predictions through the spatial plan table. Then, we propose a flow-based action prediction module to infer executable actions with coordination. Finally, we propose a spatial reasoning feedback policy to refine the spatial plan table via dual-stage replanning. Extensive experiments show that SP substantially outperforms state-of-the-art baselines, achieving over 33% improvement on Meta-World and over 25% improvement on iTHOR, demonstrating strong effectiveness across 23 embodied control tasks. We additionally evaluate SP in real-world robotic experiments to verify its practical viability. SP enhances the practicality of embodied models for robotic control applications. Code and checkpoints are maintained at https://plantpotatoonmoon.github.io/SpatialPolicy/.
1 Introduction
In recent years, hierarchical embodied models combining high-level planning and low-level action have achieved remarkable success in robotic manipulation [rt2, gr2, dp3d, spvla]. Among these approaches, vision-based methods, known as visuomotor robotic manipulation, have emerged as a particularly promising direction. By leveraging rich visual cues, these methods enable effective cross-hierarchical information transfer across diverse robotic embodiments and environments [unipi, seer, avdc, videoagent, susie, grmg]. Specifically, visuomotor robotic manipulation employs a generative model to predict trajectories of future video, which are subsequently translated into executable actions through an action prediction module.
Despite significant advances, existing visuomotor robotic manipulation frameworks lack spatial awareness abilities, failing to effectively bridge visual plans to executable actions in complex environments [seer, susie]. For instance, AVDC [avdc] lacks explicit spatial modeling in its video generation process, resulting in physically implausible predictions, such as robotic arms passing through walls. Feedback module of CLOVER [clover] enhances dynamic environment adaptation, but its inability to process spatial relationships prevents handling of occluded target scenarios [seer].
@@ 1 Introduction
To address these challenges, we propose Spatial Policy (SP), a unified framework for robotic manipulation that integrates spatial-aware video generation, action prediction, and feedback. The core idea is to unify spatial modeling and reasoning via a spatial plan table that guides video generation, enables spatially consistent action prediction, and supports feedback-driven refinement during execution.
Specifically, (1) we propose a Spatial-Conditioned Embodied Video Generation module that generates visual predictions guided by a structured spatial plan table, which encodes object-goal relations and directional cues to ensure spatial coherence in imagined futures, addressing the challenge of modeling spatially meaningful video trajectories; (2) we introduce a Flow-Based Action Prediction module that infers executable actions by capturing spatial motion dynamics between predicted frames using flow and spatial coordinates, tackling the challenge of translating visual predictions into spatially consistent actions; (3) we develop a Spatial Reasoning Feedback Policy that refines the spatial plan and video execution through dual-stage replanning based on VLM feedback and rule-based corrections, solving the challenge of real-time spatial reasoning during policy execution.
Extensive experiments show that our model surpasses existing baselines, achieving 86.7% performance on Meta-World and 59.6% on iTHOR and further proving its viability through real-world robotic experiments. The contributions of this paper are summarized as follows:
@@
  • Modular design with structured spatial modeling and reasoning. SP consists of three synergistic modules: a spatial-conditioned video generation module guided by a structured spatial plan table, a flow-based action prediction that preserves spatial consistency, and a spatial reasoning feedback policy that refines execution through real-time spatial replanning.
@@
  • Comprehensive evaluation and strong performance. We conduct extensive experiments on the MetaWorld benchmark to assess the impact of spatial integration, including ablations on video generation, action prediction, and feedback mechanisms. SP yields an absolute improvement of over 33% on Meta-World and over 25% on iTHOR and futher validate our model on real-world robotic executions, demonstrating the effectiveness of spatial reasoning in visuomotor robotic manipulation.
2 Related Work
2.1 Video Generation for Embodied Manipulation
@@
Video generation has recently been explored as a visual planning mechanism for embodied agents. AVDC [avdc] generates action-annotated video plans from raw RGB inputs using dense correspondencesenabling policy learning without action supervision. However, due to the lack of explicit spatial reasoning in both generation and action prediction, limiting precision in manipulation tasks. VideoAgent [videoagent] refines video plans via self-conditioning to reduce hallucinations and improve visual quality. However, it inherits the lack of explicit spatial reasoning from AVDC in both generation and action prediction, limiting precision in geometry-sensitive manipulation tasks. Clover [clover] introduces a closed-loop framework combining video diffusion planning and feedback control via an embedding-based error signal, achieving strong results in long-horizon tasks. However, its spatial modeling is limited to depth-based cues and lacks explicit 3D spatial planning or structured geometric reasoning, which restricts generalization across diverse spatial layout. SEER [seer] proposes a Predictive Inverse Dynamics Model (PIDM) that forecasts future visual states and infers actions end-to-end, closing the loop between vision and control. While it achieves strong performance through large-scale training, it lacks explicit spatial reasoning modules and does not leverage structured spatial plans, which may limit interpretability and adaptability in geometry-sensitive tasks. SuSIE [susie] leverages an image-editing diffusion model to generate visual subgoals for a low-level goal-conditioned policy, enabling better generalization in unstructured environments. However, its subgoal generation relies on appearance-based edits without explicit incorporation of 3D spatial relationships, limiting its precision in spatially complex manipulation scenarios.
2.2 Spatial Reasoning in Vision-Language Embodied Models
@@
Spatial understanding has been addressed in embodied AI from several perspectives. SpatialVLM [spatialvlm] and GraspCorrect [graspcorrect] leverage pretrained vision-language models to extract chain-of-thought spatial cues and determine object relations in manipulation tasks. 3D-VLA [3dvla] explicitly encodes spatial semantics by aligning 3D vision with textual instructions. SPA [spa] utilizes point cloud reconstruction for spatial alignment in manipulation and navigation tasks. Meanwhile, LLM-Planner [llmplanner] and related methods employ large language models to infer spatially structured plans in symbolic form. Despite these advances, existing methods primarily address semantic or symbolic spatial reasoning and lack visual feedback coupled with executable rollouts. Many rely on handcrafted prompts or fixed spatial priors and do not support continuous video prediction. Our approach fills this gap by embedding structured spatial plans, automatically derived from the environment geometry, into a generative video prediction framework, thereby integrating spatial reasoning and visual synthesis for action-aware planning.
@@
Figure 1: Framework Overview of Spatial Policy. Our system comprises three modules: (1) Spatial-Conditioned Embodied Video Generation, which uses a Spatial Plan Table obtained via VLM reasoning over the spatial offset between the robot and the object, to guide diffusion model in generating spatially coherent video prediction. (2) Flow-Based Action Prediction, which converts the generated video into executable actions using flow and spatial coordinates. (3) Spatial Reasoning Feedback Policy, which enables real-time correction via a dual-stage replanning strategy, combining VLM-based video judgement and policy diagnostics to refine the spatial plan table for closed-loop control.
3 Spatial Policy
We propose a visuomotor robotic manipulation framework that integrates structured spatial reasoning into the perception-to-control pipeline, enabling robots to generate spatially consistent visual plans and execute them reliably. As illustrated in Fig. 1, our system comprises three interconnected modules. First, the Spatial-Conditioned Embodied Video Generation module in Sec. 3.1 generates future video trajectories conditioned on a structured spatial plan table composed of atomic actions, directional vectors, and relative distances. Second, the Flow-Based Action Prediction module in Sec. 3.2 predicts actions aligned with the generated video using the flow and spatial coordinates. Third, the Spatial Reasoning Feedback Policy monitors spatial alignment during execution and dynamically refines both the spatial plan table and the visual predictions. Together, these modules form a closed-loop framework that maintains spatial consistency across video imagination, action prediction, and feedback-based reasoning. The full inference pipeline is provided in Appendix A.1.
3.1 Spatial-Conditioned Embodied Video Generation
@@
To generate spatially grounded visual trajectories for visuomotor robotic manipulation, we adopt a conditional diffusion framework for future video prediction. The model takes as input an initial observation frame and a textual task description, and generates a sequence of future frames representing high-level visual plans. To enhance spatial awareness, we additionally introduce the spatial plan table derived from spatial relationships as an auxiliary global condition.
@@
To capture precise spatial intent, we extract the coordinates of the robot end-effector $\mathbf{p}_{\text{ee}}\in\mathbb{R}^{3}$ and the target object $\mathbf{p}_{\text{obj}}\in\mathbb{R}^{3}$ from the simulator, and compute their relative offset as $\Delta\mathbf{p}=\mathbf{p}_{\text{obj}}-\mathbf{p}_{\text{ee}}$. This spatial offset is passed into a pretrained visual-language model (GPT-4o) to produce a structured spatial plan. The generated plan consists of sequential subgoals, each encoding an atomic action with an action type, a direction vector, and a distance scalar, as illustrated in Fig. 1. The detailed VLM inference procedure is provided in Appendix A.5.
@@
All subgoal embeddings $\{\mathbf{e}_{1},\dots,\mathbf{e}_{n}\}$ are concatenated with the task text embedding $\mathbf{z}_{\text{task}}\in\mathbb{R}^{d}$ obtained via CLIP [clip] to form the composite global condition:
@@
We follow the design of a U-Net-based denoising diffusion probabilistic model, adapted from the Imagen architecture [imagen], consistent with AVDC [avdc]. Given an initial frame $\mathbf{I}_{0}\in\mathbb{R}^{H\times W\times 3}$ and a composite condition vector $\mathcal{T}$ that incorporates both task instructions and structured spatial information, the model predicts a sequence of $T=7$ future frames $\mathbf{I}_{1:T}$, resulting in an 8-frame trajectory. The generative objective is to approximate the posterior distribution:
@@
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is sampled noise, $\beta_{t}$ is the diffusion schedule, and $\epsilon_{\theta}$ is the model’s denoising network.
3.2 Flow-Based Action Prediction
To incorporate explicit motion cues into action generation, we extend the original module by adding flow as an additional conditioning signal, while preserving the overall framework and notation style.
@@ 3.2 Flow-Based Action Prediction
We condition our policy on four types of information: (1) the current image $\mathbf{I}_{\text{cur}}$, (2) the goal image $\mathbf{I}_{\text{goal}}$ (sampled from the generated visual plan within the next $k$ steps), (3) the spatial coordinate of the robotic arm $\mathbf{p}_{\text{cur}}\in\mathbb{R}^{3}$, and (4) the dense flow field computed between consecutive frames. The flow field is computed using the classical Farnebäck method [farneback]:
$\mathbf{F}_{t}=\text{Flow}\!\left(\mathbf{I}_{t-1},\mathbf{I}_{t}\right),$ (5)
and then encoded using a CNN-based encoder. All inputs are encoded into compact feature vectors through their respective encoders (ResNet for images and an MLP for spatial coordinates). The modality-specific features are subsequently concatenated to form the global condition vector used by the diffusion policy.
Diffusion Policy-based Action Prediction Module
We formulate action prediction as a conditional generative process. Specifically, we adopt a U-net-based diffusion policy architecture to model the conditional distribution over action trajectories:
$p_{\theta}(\mathbf{a}_{0:H-1}\mid\mathbf{I}_{\text{cur}},\mathbf{I}_{\text{goal}},\mathbf{p}_{\text{cur}},\mathbf{F}_{t})$ (6)
@@
where $\mathbf{c}$ denotes the concatenated global condition vector including all encoded modalities, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise, and $t$ is the diffusion timestep. $\epsilon_{\theta}$ is the learned denoising network. At inference time, the model predicts the entire action trajectory in a denoising pass, and the resulting sequence is executed step-by-step in the environment. The frame–switching mechanism used during inference is described in Appendix A.4.
3.3 Spatial Reasoning Feedback Policy